# FOSP: Fine-tuning Offline Safe Policy through World Models

### v1

@@ A.1 Derivation of the Final Optimization Problem
$\max_{\pi}\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi}[A(s,a)]\;\;\text{s.t.}\;D(\pi_{\psi}||\pi_{b})\leq\epsilon.$
@@
$L(\pi,\lambda,\mu)=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi}[A({\bm{s}},{\bm{a}})]-\lambda(D(\pi_{\psi}||\pi_{b})-\epsilon),$
@@
$\frac{\partial L}{\partial\pi}=A(s,a)-\lambda\log\pi_{b}({\bm{a}}|{\bm{s}})+\lambda\log\pi({\bm{a}}|{\bm{s}})+\lambda=0$
@@
Therefore, we have a solution to our optimization objective (13). The solved $\pi_{\psi}$ can be explicitly expressed by Eq.(15).
@@ Appendix B Implementation Details
The TD($\lambda$) return can be estimated as follows:
$R_{t}^{H}=V_{\varphi}({\bm{s}}_{t}^{H}),\;R_{t}^{h}={\bm{r}}_{t}^{h}+\gamma((1-\lambda)V_{\varphi}({\bm{s}}_{t}^{h})+\lambda R_{t}^{h+1})$ (22)
where $H$ is the horizon of the model and $h\leq H$.
To implement the reachability estimation function $u_{\xi}$, we use the theorem introduced by [51] (Theorem 1). The REF function can be trained as follows:
$u({\bm{s}}_{t})=\max\{\mathds{1}_{s\in\mathcal{S}_{f}},\gamma_{u}u({\bm{s}}_{t+1})\},$ (23)
where $\gamma_{u}$ is a discount parameter $10<<\gamma_{u}<$ 1 to ensure convergence of $u({\bm{s}})$. Meanwhile, we ensure that its learning rate is greater than that of the critic and policy, but less than that of the Lagrangian multiplier updates.
@@
1:  Initialize neural network parameters $\theta,\psi,\phi_{i},\varphi_{i},\xi$.$(i=1,2)$
@@
3:     Sample a batch of trajectories $B\sim\mathcal{D}$.
@@
5:     Update $\theta$ using Eq. (D).
@@
8:     Update $V_{\varphi_{1}},V_{\varphi_{2}}$ using Eq. (6).
9:     Update $Q_{\phi_{1}},Q_{\phi_{2}}$ using Eq. (8).
10:     Update ${\pi}_{\psi}$ using Eq. (15).
11:     Update $u_{\xi}$ using Eq. (23).
@@
14:     Switch a policy $\pi$ according to Eq.(9) and rollout the policy in the environment for an episode to collect a new trajectory $\tau$.
@@
18:        Update $\theta$ using Eq. (D).
@@
21:        Update $V_{\varphi_{1}},V_{\varphi_{2}}$ using Eq. (6).
22:        Update $Q_{\phi_{1}},Q_{\phi_{2}}$ using Eq. (8).
23:        Update ${\pi}_{\psi^{\prime}}$ using Eq. (17).
24:        Update $u_{\xi}$ using Eq. (23).
Appendix C Experimental Details
C.1 Real World Experiments
The experiments verify the effectiveness of the algorithm by controlling a robotic arm to avoid obstacles and complete a reach task to the target objection, called SafeReach. As shown in Figure 4(c), the main hardware setup consists of three parts: a Franka Panda robotic arm, an experimental platform for placing obstacles and target objections, and an Intel RealSense D435 camera for perception. On the experimental platform, we used white lines to designate a 25cm×$\times$25cm operational area as the task space for the robotic arm.
@@ C.1 Real World Experiments
For the task of reaching, the robot needs to consider the image information acquired by the camera along with the current joint posture information as planning conditions to determine the movement decision at the next step. As shown in Figure 4, the robotic arm needs to navigate within the designated movement space, bypassing the obstacles which are represented by the green rectangular blocks, to touch the predefined target which is represented by the red square block.
@@
To train the model for the SafeReach task, 200 motion trajectories were collected as a dataset by teleoperating the robotic arm using a 3dconnexion SpaceMouce. The dataset consists of the following: 120 trajectories where the robotic arm did not touch any obstacles and successfully reached the target object; 30 trajectories where it did not touch any obstacles but also did not successfully reach the target; 40 trajectories where it touched obstacles but successfully reached the target; 10 trajectories where it touched obstacles and did not successfully reach the target. Each trajectory contains 80 sets of data points at different time steps. Each data point includes: state: a 64$\times$64 RGB image captured by the camera, a 6-dimensional pose list of the robotic arm; action: a 6-dimensional list recording the changes in the robotic arm’s pose; cost: 1-dimensional cost index indicating whether the robotic arm touched an obstacle; reward 1-dimensional reward index indicating whether the object reached the target object. The labeling methods for the cost and reward index are shown in Figure 6. The cost is set to 1 when the robotic arm collides with an obstacle and it is set to 0 otherwise. The reward is set to 1 when the object touches the target and it is set to 0 otherwise. We defined 3 placement points for target objects and 6 placement points for obstacles in the scenario. By restricting the placement positions of the target objects and obstacles, we ensure that the model can converge more efficiently on a limited dataset.
@@
To validate the generalization of the algorithm, we design a series of transfer experiments. As illustrated in the left side of Figure 5, the original task involves 2 green rectangular obstacles and 1 red cylindrical target objection. We design three different transfer tasks, depicted on the right side of Figure 5. The tasks include altering the shape of the obstacles, changing the shape of the target object, and increasing the number of obstacles. These experiments aim to verify the algorithm’s generalization in avoiding obstacles of various shapes and quantities, as well as in reaching target objects of different shapes.
C.2 Simulation Environment
@@
C.3 Simulation Dataset
We collect simulation data from three policies: random policy, safe policy and unsafe policy. Each policy collected 200 trajectories for each task, which were then combined into the final dataset. This approach ensures that our dataset uniformly includes all types of data: low cost with low return, low cost with moderate return, high cost with low return, and high cost with high return. Thus, it can make a better trade-off in training different offline policies. The safe policy is trained by SafeDreamer[8] and the unsafe policy is trained by DreamerV3[5] for 1.5M steps on each task. It shows that unsafe policies tend to collect data with higher returns but significant cost uncertainty in these scenarios.
Figure 7: The dataset distribution on five simulation tasks. We plot the reward return and the cost return of every data trajectory.
Appendix D Baselines
Figure 8: Online experimental results. Comparing FOSP to baselines across five image-based safety tasks in online fine-tuning. The results for all three algorithms are obtained after fine-tuning for 750,000 steps. As we can see, FOSP can maintain safety and achieve better performance during fine-tuning.
@@ Appendix D Baselines
DreamerV3[5] is a model-based reinforcement learning method that outperforms specialized methods across over 150 diverse tasks. It has a stable performance without adjusting hyperparameters in exploring farsighted strategies from pixels and sparse rewards in an open world. However, it overlooks the safety considerations in the environment, which brings high costs in safety-critical tasks.
It uses RSSM as the world model, and the loss function is similar to Eq.(D) but lacks the cost head.
$\mathcal{L}^{\text{model}}(\theta)=$ $\operatorname*{\mathbb{E}}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}-\ln p_{\theta}({\bm{x}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{r}}_{t}\mid{\bm{s}}_{t})+\mathbb{D}_{KL}[sg(q_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t},{\bm{x}}_{t}))||p_{\theta}^{i_{t}}({\bm{z}}_{t}|{\bm{h}}_{t})]+$
@@
In the offline pretraining phase, we removed the online planning component and used it only for online fine-tuning. To achieve robust performance in the offline phase, we train a latent dynamics model ensemble and use the uncertainty estimation approach[14]:
@@
SafeDreamer[8] incorporates Lagrangian-based methods into world model planning processes and achieves nearly zero-cost performance on various tasks. It performs well in high-dimensional vision-only input safety-critical tasks, surpassing the prior works[7, 40] and balancing performance and safety. The framework uses the same world models based on DreamerV3[5]. We use the BSRP-Lag version of the model, which utilizes the Lagrangian method in background safety-reward planning that avoids online planning. And this version of SafeDreamer also achieves the best performance of the three versions.
It adds a cost head into world models like our method and uses the same loss Eq.(D) in model training. It constructs a cost critic $V_{\phi}^{c}(R_{t}|{\bm{s}}_{t})$ like $V_{\phi}(R_{t}|{\bm{s}}_{t})$ and utilize Augmented Lagrangian method to update the actor:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}R^{\lambda}(s_{t})+\eta\mathrm{H}\left[\pi_{\theta}\left(a_{t}\mid s_{t}\right)\right]-\Psi\left(C^{\lambda}(s_{t}),\lambda_{p}^{k},\mu^{k}\right)$
$\Psi\left(C^{\lambda}(s_{t}),\lambda_{p}^{k},\mu^{k}\right),\lambda_{p}^{k+1}=\begin{cases}\lambda_{p}^{k}\Delta+\frac{\mu^{k}}{2}\Delta^{2},\lambda_{p}^{k}+\mu^{k}\Delta&\text{ if }\lambda_{p}^{k}+\mu^{k}\Delta\geq 0\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{2\mu^{k}},0&\text{ otherwise }\end{cases}$
where $\delta=C^{\lambda}(s_{t})-b$.
Appendix E Hyperparameters
The experiments for FOSP were conducted in a Python 3.10 environment with JAX 0.4.26. Our setup included CUDA version 12.1, running on Ubuntu 20.04. The hardware used comprised of four GeForce RTX 4090 GPUs and an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz CPU.
@@ Appendix E Hyperparameters
General Number of MLP layers 5
Number of MLP layer units 512
Action repeat 4
Appendix F Additional Experimental Results

### v2

@@ A.2 Extraction of the Optimal Policy
$\max_{\pi}\mathbb{E}_{{\bm{a}}\sim\pi}[A({\bm{s}},{\bm{a}})]\;\;\text{s.t.}\;D_{\mathrm{KL}}(\pi_{\psi}||\pi_{b})\leq\epsilon.$
@@
$L(\pi_{\psi},\lambda,\mu)=\mathbb{E}_{{\bm{a}}\sim\pi}[A({\bm{s}},{\bm{a}})]-\lambda(D_{\mathrm{KL}}(\pi_{\psi}||\pi_{b})-\epsilon),$
@@
$\frac{\partial L}{\partial\pi_{\psi}}=A({\bm{s}},{\bm{a}})-\lambda\log\pi_{b}({\bm{a}}|{\bm{s}})+\lambda\log\pi_{\psi}({\bm{a}}|{\bm{s}})=0.$
@@
Therefore, we have a solution to equation 16. The solved $\pi_{\psi}$ can be explicitly expressed by equation 20.
@@ Appendix B Implementation Details
To implement the REF $u_{\xi}$, we use the theorem introduced by Ganai et al. (2024) (Theorem 1). The REF can be trained as follows:
$u({\bm{s}}_{t})=\max\{\mathds{1}\{s\in\mathcal{S}_{f}\},\gamma_{u}u({\bm{s}}_{t+1})\},$ (30)
where $\gamma_{u}$ is a discount parameter $10\ll\gamma_{u}<$ 1 to ensure convergence of $u({\bm{s}})$. Meanwhile, we ensure that its learning rate is greater than that of the critic and policy, but less than that of the Lagrangian multiplier updates.
Appendix C Pseudo Code
@@
1:  Initialize neural network parameters $\theta,\psi,\phi_{k},\varphi_{k},\xi$. $(k=1,2)$
@@
3:     Sample a batch of trajectories $\mathcal{B}\sim\mathcal{D}$.
@@
5:     Update $\theta$ using equation F.
@@
8:     Update $V_{\varphi_{1}},V_{\varphi_{2}}^{c}$ using equation 7.
9:     Update $Q_{\phi_{1}},Q_{\phi_{2}}^{c}$ using equation 11.
10:     Update ${\pi}_{\psi}$ using equation 20.
11:     Update $u_{\xi}$ using equation 30.
@@
14:     Switch a policy $\pi$ according to equation 21 and rollout the policy in the environment for an episode to collect a new trajectory $\tau$.
@@
18:        Update $\theta$ using equation F.
@@
21:        Update $V_{\varphi_{1}},V_{\varphi_{2}}^{c}$ using equation 7.
22:        Update $Q_{\phi_{1}},Q_{\phi_{2}}^{c}$ using equation 11.
23:        Update ${\pi}_{\psi^{\prime}}$ using equation 23.
24:        Update $u_{\xi}$ using equation 30.
Appendix D Experimental Details
D.1 Real World Experiments
The experiments verify the effectiveness of the algorithm by controlling a robotic arm to avoid obstacles and complete a Reach task to the target object, called SafeReach. As shown in Figure 8(c), the main hardware setup consists of three parts: a Franka Panda robotic arm, an experimental platform for placing obstacles and target objects, and an Intel RealSense D435 camera for perception. On the experimental platform, we used white lines to designate a 25cm×$\times$25cm operational area as the task space for the robotic arm.
@@ D.1 Real World Experiments
For the task of reaching, the robot receives the image information acquired by the camera along with the current joint posture information as planning conditions to determine the movement decision at the next step. As shown in Figure 8, the robotic arm needs to navigate within the designated movement space, bypassing the obstacles which are represented by the green rectangular blocks, to touch the predefined target which is represented by the red square block.
@@
To train the model for the SafeReach task, 200 motion trajectories were collected as a dataset by teleoperating the robotic arm using a 3Dconnexion SpaceMouse. The dataset consists of 120 trajectories where the robotic arm did not touch any obstacles and successfully reached the target object; 30 trajectories where it did not touch any obstacles but also did not successfully reach the target; 40 trajectories where it touched obstacles but successfully reached the target; 10 trajectories where it touched obstacles and did not successfully reach the target. Each trajectory contains 80 data points at different time steps. Each data point includes state: a 64$\times$64 RGB image captured by the camera, a 6-dimensional pose of the robotic arm; action: a 6-dimensional variation recording the changes in the robotic arm’s pose; cost: 1-dimensional cost index indicating whether the robotic arm touched an obstacle; reward: 1-dimensional reward index indicating whether the object reached the target object. The labeling methods for the cost and reward index are shown in Figure 9. The cost is assigned a value of 1 when the robotic arm collides with an obstacle, and 0 otherwise. Similarly, the reward is set to 1 when the object reaches the target and 0 in all other cases. We defined 3 placement points for target objects and 6 placement points for obstacles in the scenario. By restricting the placement positions of the target objects and obstacles, we ensure that the model can converge more efficiently on a limited dataset.
@@
To validate the generalization of the algorithm, we design a series of transfer experiments. As illustrated in the upper side of Figure 7, the original task involves 2 green rectangular obstacles and 1 red cylindrical target object. We design three different transfer tasks, depicted on the lower side of Figure 7. The tasks include altering the shape of the obstacles, changing the shape of the target object, and increasing the number of obstacles. These experiments aim to verify the algorithm’s generalization in avoiding obstacles of various shapes and quantities, as well as in reaching target objects of different shapes.
D.2 Simulation Environment
Metrics
In the simulation environment, we use three metrics to measure the performance:
  • Reward $J_{r}$ is the average episodic sum of rewards.
    $J_{r}=\frac{1}{E}\sum^{E}_{i=1}\sum^{T_{ep}}_{t=1}r_{t,i},$
  • Cost $J_{c}$ is the average episodic sum of costs.
    $J_{c}=\frac{1}{E}\sum^{E}_{i=1}\sum^{T_{ep}}_{t=1}c_{t,i},$
  • Cost Regret $\rho_{c}$ means the average cost over the entirety of training.
    $\rho_{c}=\frac{\sum^{T}_{t=1}c_{t}}{T},$
    where E is the number of episodes, $T_{ep}$ is the timestep of episodes and $T$ is the total timestep.
@@ D.2 Simulation Environment
SafetyFading
This environment is similar to the SafetyGoal, where the agent needs to reach a goal while avoiding obstacles. However, over time, the goal and obstacles gradually become undetectable. The agent needs to gather as much information about the environment as possible from the initial stage and form a memory of the goal and obstacle locations to complete the task according to its initial plan. This is undoubtedly more challenging than SafetyGoal and is even harder to accomplish for visual-only input agents.
D.3 Simulation Dataset
We collect the simulation standard dataset from three policies: random policy, safe policy and unsafe policy. Each policy collected 200 trajectories for each task, which were then combined into the final dataset. This approach ensures that our dataset uniformly includes all types of data: low cost with low return, low cost with moderate return, high cost with low return, and high cost with high return. Thus, it can make a better trade-off in training different offline policies. The safe policy is trained by SafeDreamer (Huang et al., 2023) and the unsafe policy is trained by DreamerV3 (Hafner et al., 2023) for 1.5M steps on each task. The standard dataset distributions on five simulation tasks are illustrated in Figure 10. They show that unsafe policies tend to collect data with higher returns but significant cost uncertainty in these scenarios.
Figure 10: The dataset distribution on five simulation tasks. We plot the reward return and the cost return of every data trajectory. Gray dots represent random data, yellow dots represent unsafe data, and blue dots represent safe data.
Appendix E Details of Real-world Experiments
We illustrate the details of real-world experiments in Figure 11. The FOSP is first pretrained on an offline dataset for 0.5 million steps and deployed in the real world to execute the unseen tasks. In each task, we recorded the robot’s trails before and after fine-tuning over 40 trials. As we can see, the robot’s ability to safely complete unseen tasks has significantly improved after fine-tuning. It shows the strength of fine-tuning on different tasks.
We also compare our method with SafeDreamer on these tasks after fine-tuning. Both methods are pretrained for 500,000 steps and undergo 40 fine-tuning iterations. As depicted in Figure 12, SafeDreamer sometimes violates constraints even when the task is completed. In contrast, FOSP can learn with zero constraint violations and ultimately reach the goal.
Figure 11: Deployment on different real-world unseen tasks. The trails from top to bottom show the original offline training task and three unseen fine-tuning tasks, with each fine-tuning task displaying both before and after fine-tuning trails. The green frames indicate successful task completion, while the orange frames represent failures.
Figure 12: Compare with baseline on different real-world unseen tasks. In each task, we compare our method with the fine-tuned SafeDreamer (baseline). The green frames indicate successful goal reaching, the red frames indicate constraint violations and the orange frames represent failures.
Appendix F Baselines
@@
DreamerV3 (Hafner et al., 2023) is a model-based reinforcement learning method that outperforms specialized methods across over 150 diverse tasks. It has a stable performance without adjusting hyperparameters in exploring farsighted policies from pixels and sparse rewards in an open world. However, it overlooks the safety considerations in the environment, which brings high costs in safety-critical tasks.
It uses RSSM as the world model, and the loss function is similar to equation F but lacks the cost head.
$\mathcal{L}^{\text{model}}(\theta)=$ $\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}-\ln p_{\theta}({\bm{x}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{r}}_{t}\mid{\bm{s}}_{t})+\mathbb{D}_{KL}[sg(q_{\theta}({\bm{z}}_{t}|{\bm{h}}_{t},{\bm{x}}_{t}))||p_{\theta}^{i_{t}}({\bm{z}}_{t}|{\bm{h}}_{t})]+$
@@
In the offline pretraining phase, we removed the online planning component and used it only for online fine-tuning. To achieve robust performance in the offline phase, we train a latent dynamics model ensemble and use the uncertainty estimation approach (Yu et al., 2020):
@@
SafeDreamer (Huang et al., 2023) incorporates Lagrangian-based methods into world model planning processes and achieves nearly zero cost performance on various tasks. It performs well in high-dimensional vision-only input safety-critical tasks, surpassing the prior works (As et al., 2022; Hogewind et al., 2022) and balancing performance and safety. The framework uses the same world models based on DreamerV3 (Hafner et al., 2023). We use the BSRP-Lag version of the model, which utilizes the Lagrangian method in background safety-reward planning that avoids online planning. In the offline setting, we train it by providing the offline dataset and removing the parts involving interaction with the environment. This version of SafeDreamer also achieves the best performance of the three versions.
It adds a cost head into world models like our method and uses the same loss equation F in model training. It constructs a cost critic $V_{\phi}^{c}(R_{t}|{\bm{s}}_{t})$ like $V_{\phi}(R_{t}|{\bm{s}}_{t})$ and utilize Augmented Lagrangian method to update the actor:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}R^{\lambda}({\bm{s}}_{t})+\eta\mathrm{H}\left[\pi_{\theta}\left({\bm{a}}_{t}\mid{\bm{s}}_{t}\right)\right]-\Psi\left(C^{\lambda}({\bm{s}}_{t}),\lambda_{p}^{k},\mu^{k}\right),$
$\Psi\left(C^{\lambda}({\bm{s}}_{t}),\lambda_{p}^{k},\mu^{k}\right),\lambda_{p}^{k+1}=\begin{cases}\lambda_{p}^{k}\Delta+\frac{\mu^{k}}{4}\Delta^{2},\lambda_{p}^{k}+\frac{\mu^{k}}{2}\Delta,&\text{ if }\lambda_{p}^{k}+\frac{\mu^{k}}{2}\Delta\geq 0,\\ -\frac{\left(\lambda_{p}^{k}\right)^{2}}{\mu^{k}},0,&\text{ otherwise, }\end{cases}$
where $\delta=C^{\lambda}({\bm{s}}_{t})-b$.
Recovery RL
Recovery RL (Thananjeyan et al., 2021) leverages offline data to learn the constraint violation zones before learning and uses two policies to separate the goal of enhancing performance and satisfying the constraints. It achieves nearly zero cost in uncertain environments where safety limits exploration. We use the model-based version of this baseline. First, we train cost Q-value by the following MSE loss:
$\mathcal{L}(\phi)=(Q_{\phi}^{c}({\bm{s}}_{t},{\bm{a}}_{t})-({\bm{c}}_{t}+(1-{\bm{c}}_{t})\gamma_{c}\mathbb{E}_{\pi}[Q_{\phi}^{c}({\bm{s}}_{t+1},{\bm{a}}_{t+1})]))^{2}.$
Then we select actions from the safe set and the recovery set:
${\bm{a}}_{t}=\begin{cases}{\bm{a}}_{t}^{\pi_{\mathrm{task}}},&\text{ if }({\bm{s}}_{t},{\bm{a}}_{t})\in\{({\bm{s}},{\bm{a}})\in\mathcal{S}\times\mathcal{A}:Q^{c}_{\phi}({\bm{s}},{\bm{a}})\leq\epsilon_{c}\},\\ {\bm{a}}_{t}^{\pi_{\mathrm{recovery}}},&\text{ if }({\bm{s}}_{t},{\bm{a}}_{t})\in\{({\bm{s}},{\bm{a}})\in\mathcal{S}\times\mathcal{A}:Q^{c}_{\phi}({\bm{s}},{\bm{a}})>\epsilon_{c}\},\end{cases}$
where $\epsilon_{c}$ is a threshold. We follow the model predictive control (MPC) as a learned dynamic model $f_{\theta}$ and use a VAE-based model to capture the high-dimensional information. And we utilize SAC (Haarnoja et al., 2018) to learn $\pi_{\mathrm{task}}$.
PPO-Lagrangian
PPO-Lagrangian uses the objective of clipped PPO (Schulman et al., 2017) to optimize:
$\mathcal{L}(\theta)_{ppo}=\min(\frac{\pi_{\theta}({\bm{a}}|{\bm{s}})}{\pi_{\theta_{k}}({\bm{a}}|{\bm{s}})}A_{r}^{\pi_{\theta_{k}}}({\bm{s}},{\bm{a}}),\text{clip}(\frac{\pi_{\theta}({\bm{a}}|{\bm{s}})}{\pi_{\theta_{k}}({\bm{a}}|{\bm{s}})},1-\epsilon_{\text{clip}},1+\epsilon_{\text{clip}})A_{r}^{\pi_{\theta_{k}}}({\bm{s}},{\bm{a}})),$
We use the PID Lagrangian (Stooke et al., 2020) method and obtain the loss of the PPO-Lagrangian:
$\mathcal{L}(\theta)_{ppol}=\frac{1}{1+\lambda}(\mathcal{L}(\theta)_{ppo}-\lambda A_{c}^{\pi_{\theta_{k}}}({\bm{s}},{\bm{a}})).$
CPO
CPO (Achiam et al., 2017) uses a local policy search combined with trust region recovery to ensure that single-step policy updates follow a direction that does not violate the constraints. It introduces this form to optimize the problem:
$\theta^{*}=\theta_{k}-\sqrt{\frac{2\delta}{b^{T}H^{-1}b}}H^{-1}b,$
where H is the e Hessian of KL-divergence.
@@
REF Learning rate $l_{r}$ $5\cdot 10^{-5}$
General Number of MLP layers $N_{\text{MLP}}$ 5
Number of MLP layer units $N_{\text{units}}$ 512
Action repeat $n_{\text{repeat}}$ 4
Appendix G Hyperparameters
The experiments for FOSP were conducted in a Python 3.10 environment with JAX 0.4.26. Our setup included CUDA version 12.1, running on Ubuntu 20.04. The hardware used comprised four GeForce RTX 4090 GPUs and an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz. And the experiments’ hyperparameters setting is shown in Table 4.
Figure 13: Offline experimental results. Comparing FOSP to baselines across five image-based safety tasks. The results for all three algorithms are obtained after training for 1 million steps. Reward: averaged episode reward return. Cost: averaged episode cost return. Cost Regret: averaged cost value throughout the training phase. As we can see, FOSP can maintain safety and achieve better performance during offline training.
Figure 14: Online experimental results without DreamerV3. These results are the same as Figure 3. We omit the curves of some baselines to clearly illustrate the results of other main methods.
Appendix H Additional Experimental Results
H.1 Offline Training Results